OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?

OpenStudy (anonymous):

first find the slope: m=(y1-y2)/(x1-x2) then stick the slope in point slope form: y-y1 = m(x-x1)